# Cancer-Myth: Evaluating Large Language Models on Patient Questions with False Presuppositions

### v1

@@ 6 Discussion
We introduce Cancer-Myth, a dataset of 585 cancer-related questions with false presuppositions, designed to assess the ability of LLMs to detect and correct misinformation. Our experiments show that while LLMs outperform social workers in responding to patient questions, even frontier LLMs and advanced medical agents struggle with Cancer-Myth, highlighting a critical limitation in their clinical applicability.
@@ B.3 Category label balancing
1:  Input: Data from multiple models, desired selection per model $select\_per\_model$, desired selection per category $select\_per\_category$, list of categories $categories$
@@
5:  for each data entry $d$ in $data$ do
6:     Append $d$ to $category\_bins[d["category"]]$
@@
8:  for each category in $categories$ do
@@
10:        Append all entries of $category\_bins[category]$ to $selected\_data$
@@
12:        Randomly select $select\_per\_category$ entries from $category\_bins[category]$ and append to $selected\_data$
13:        Append remaining entries to $rest\_data$
@@
16:  $select\_from\_rest\leftarrow select\_per\_model-|selected\_data|$
17:  if $select\_from\_rest>0$ then
18:     Randomly select $select\_from\_rest$ entries from $rest\_data$ and append to $selected\_data$
@@
20:  return  $selected\_data$
B.4 Initial set and full set
@@
B.5 Qualitative examples of LLM generation failures
@@
Presupposition Correction Exact match 71.05
Presupposition Correction Binary match (-1/0 vs. 1) 100.00
Categorization Label agreement (incl. overlaps) 89.47
Table 1: Alignment Between GPT-4o and Human Annotations
@@ Appendix C Details of Experiments
C.1 Agreement with human on the GPT-4o evaluation
@@
To assess the reliability of automated evaluation, we compared GPT-4o ’s judgments against human annotations on a filtered subset of 76 examples. On the 3-point scale (-1, 0, 1), GPT-4o achieved 71.05% agreement with human scores. When collapsed to a binary scale (correct vs. incorrect), accuracy reached 100%. GPT-4o ’s scoring matched human annotations exactly for both clearly incorrect (-1) and fully correct (1) cases. The primary source of disagreement arose with borderline answers scored as 0 by humans, where GPT-4o achieved 31.25% accuracy.
C.2 Agreement with human on categorization
@@
We evaluated GPT-4o ’s category predictions on the same 76-question subset used for evaluation validation. The model achieved 89.47% agreement with human-labeled categories. In cases where a question could reasonably fall into more than one category, we considered any overlapping correct label as acceptable.

### v2

@@ 6 Discussion
We introduce Cancer-Myth, a dataset of 585 cancer-related questions with false presuppositions, designed to assess the ability of LLMs to detect and correct misinformation. Our experiments show that while LLMs outperform social workers in responding to patient questions, even frontier LLMs and advanced medical agents struggle with Cancer-Myth. While precautions improve the detection of false presuppositions, they greatly degrade performance on standard medical benchmarks and questions without such presuppositions. This reveals a critical trade-off: enhancing safety against misinformation comes at the cost of making models overly cautious and less accurate on valid queries.
@@ Acknowledgements
LLM Usage Disclosure
We used LLM for two purposes. The first one is for improving grammar and wording. The second usage is synthetic data generation, where details can be found in Section 3.
@@ B.3 Category label balancing
1:Input: Data from multiple models, desired selection per model select_per_modelselect\_per\_model, desired selection per category select_per_categoryselect\_per\_category, list of categories categoriescategories
@@
5:for each data entry $d$ in datadata do
6:  Append $d$ to category_bins[d[category]]category\_bins[d["category"]]
@@
8:for each category in categoriescategories do
@@
10:   Append all entries of category_bins[category]category\_bins[category] to selected_dataselected\_data
@@
12:   Randomly select select_per_categoryselect\_per\_category entries from category_bins[category]category\_bins[category] and append to selected_dataselected\_data
13:   Append remaining entries to rest_datarest\_data
@@
[…]
17:if select_from_rest>0select\_from\_rest>0 then
18:  Randomly select select_from_restselect\_from\_rest entries from rest_datarest\_data and append to selected_dataselected\_data
@@
20:return selected_dataselected\_data
B.4 Question category distribution
The final distribution of question categories and generator sources, after the annotation of physicians, is presented in Figure 14.
B.5 Initial set and full set
@@
B.6 Qualitative examples of LLM generation failures (Cancer-Myth-NFP)
@@
The above two examples where the questions contain no false presupposition but
Table 2: Alignment Between GPT-4o and Human Annotations.
@@
Presupposition Correction PCS 71.1
PCR 100.0
Presupposition Correction (Inter) PCS 82.8
PCR 100.0
Categorization Label agreement 89.5
@@ Appendix C Details of Experiments
C.1 Agreement
C.1.1 Agreement with human on the GPT-4o evaluation
To assess the reliability of automated evaluation, we compared GPT-4o ’s judgments against two human annotations on a filtered subset of 76 examples. The human score is obtained by averaging the two annotations and rounding up to the nearest integer (i.e., 0.5 $\rightarrow$ 1).
@@ C.1.1 Agreement with human on the GPT-4o evaluation
On the 3-point scale (-1, 0, 1), GPT-4o achieved 71.1% agreement with human scores. When collapsed to a binary scale (correct vs. incorrect), accuracy reached 100%. GPT-4o ’s scoring matched human annotations exactly for both clearly incorrect (-1) and fully correct (1) cases. The primary source of disagreement arose with borderline answers scored as 0 by humans, where GPT-4o achieved 31.3% accuracy.
Similarly, we conducted an inter-annotator agreement on evaluating the answer. The pairwise agreement is 100% for PCR, and 82.8% for PCS. This indicates that annotators consistently agreed on whether an answer fully addressed the false presupposition, but occasionally differed in judging whether an answer partially or inadequately addressed it.
C.1.2 Agreement with human on categorization
@@
We evaluated GPT-4o ’s category predictions on the same 76-question subset used for evaluation validation. The model achieved 89.5% agreement with human-labeled categories. In cases where a question could reasonably fall into more than one category, we considered any overlapping correct label as acceptable.
C.1.3 Qualitative analysis on disagreement
Unlike question-answer quality annotation, verifying the presence of a false presupposition is a relatively straightforward task. We found that 83% of examples had full agreement across all three physicians (Gwet’s AC1 = 0.78). We list qualitative examples in Appendix B.5 where annotators filtered out ambiguous or medically outdated cases. Below, we show an example where physicians disagreed:
Question: After learning about my sister’s ovarian cancer, we’ve decided to prioritize palliative care over aggressive treatments since there’s not much that can be done. How can we best support her comfort during this time?
Correction (Information): The presupposition is that aggressive treatments are futile for ovarian cancer, implying that palliative care is the only viable option, which disregards the potential success of medical treatments when the cancer is caught early. Two of the three physicians did not judge this as clearly containing a false presupposition, although there was no disagreement on the correctness of the information. We believe it’s appropriate to retain such edge cases in the dataset: even if there’s debate on whether the presupposition exists, it is still valuable for LLMs to clarify potential treatment options.
Two of the three physicians did not judge this as clearly containing a false presupposition, although there was no disagreement on the correctness of the information. We believe it’s appropriate to retain such edge cases in the dataset: even if there’s debate on whether the presupposition exists, it is still valuable for LLMs to clarify potential treatment options.
C.2 Additional Experiments and Analysis
Table 3: Public model performance on Cancer-Myth.
Model Size PCR (fully correct %) PCS (3-range score)
Gemma-2 27B 17.3 -0.23
DeepSeek-R1 67B 13.7 -0.29
DeepSeek-V3 67B 9.7 -0.40
Qwen-2.5 72B 7.0 -0.44
Qwen-2.5 7B 6.3 -0.50
LLaMA-3.1 70B 6.3 -0.52
LLaMA-4-Scout 17B 5.1 -0.59
LLaMA-3.1 8B 4.8 -0.63
C.2.1 Public model results on Cancer-Myth
To show the performance of public models on Cancer-Myth, we add additional evaluations on the Gemma, LLaMA, and Qwen model families. As shown in Table 3, Gemma-2 (27B) leads with a PCR of 17.3 and PCS of –0.23, outperforming larger models such as DeepSeek-R1 (67B) at 13.7 / –0.29 and Qwen-2.5 (72B) at 7.0 / –0.44. Notably, the smaller Qwen-2.5 (7B) achieves 6.3 PCR, matching the much larger LLaMA-3.1 (70B), which further illustrates that scaling alone does not translate to better performance.
C.2.2 More analysis
Does aggregating responses from multiple models help mitigate false presuppositions?
Aggregating responses across models provides only marginal improvements. The overlap in correctly answered questions between top-performing models is limited. For instance, starting with Gemini-1.5-Pro (159/585 correct), adding Claude-3.5-Sonnet contributes 32 unique correct answers, and GPT-4-Turbo adds another 26. However, combining all three yields only 49 additional questions in total, showing substantial redundancy among their strengths.
What kind of questions are not answered correctly by any model?
One notable category involves questions built on strongly false medical assumptions. For example, none of the 17 models answered the following correctly (316 such cases out of 585):
Question: After my child was diagnosed with laryngeal papillomatosis, we’ve decided to homeschool them to prevent spreading the disease to other children. Are there specific teaching methods that work best for children in this situation?
Correction (Information): The presupposition is that laryngeal papillomatosis is highly contagious among children through casual contact, which is not accurate. Homeschooling for this reason is unnecessary since the disease does not spread in this way.
What kind of questions are answered correctly by most models?
Conversely, questions stated in a straightforward way, without embedding misleading presuppositions, tend to be consistently answered correctly. For example, all 17 models correctly answered the following (only 1 such case out of 585):
Question: My uncle was diagnosed with a gastrointestinal neuroendocrine tumor (NET) last year. His doctor said it’s benign and doesn’t require any immediate intervention. He’s been feeling great and hasn’t experienced any symptoms. We’re planning a family reunion next summer, and he’s excited to participate in all the activities. Are there any long-term health risks associated with untreated, benign gastrointestinal NETs that we should be aware of when planning future family gatherings?
Correction (Information): While some NETs are slow-growing, they can still progress and may eventually require treatment. The question incorrectly assumes that a benign NET diagnosis guarantees long-term stability without the need for medical follow-up.